A jar contains 9 red marbles numbered 1 to 9 and 3 blue marbles numbered 1 to 3. A marble is drawn at random from the jar. Find the probability that the marble is red or even-numbered?


1
Expert's answer
2022-05-10T18:31:03-0400

There are "9+3=12" marbles.

There are "9" red marbles. There is one blue even-numbered marble (numbered 2).

So there are "9+1" favorable outcomes and there are 12 total outcomes.

Then he probability that the marble is red or even-numbered is


"P(R\cup even)=\dfrac{9+1}{9+3}=\dfrac{5}{6}"